OpenStudy (anonymous):

suppose tanx =(√2)/4 , cos x = -(2√2)/3 find sinx cscx sec x and cot x

OpenStudy (anonymous):

tanx=sinx/cosx so tanx*cosx = sinx cscx=1/sinx secx=1/cosx cot=1/tanx

OpenStudy (anonymous):

sinx = -1/3 cscx = -3 \[\sec x = -3/2\sqrt{2}\] \[\cot x = 4/\sqrt{2}\]